A voltage source of 10 V is connected to a series RC circuit where C=3.0μF and R= 2.0×10^6 Ω. Find the amount of time required for the capacitor to reach 95% of its maximum voltage.
Answer #1

Time constant, T = RC

T = 2 x 10^6 x 3 x 10^-6 = 6 sec

Charge, Q = Qo (1 - e^(-t/T))

0.95 = 1 - e^(-t/T)

t = - T ln(1 - 0.95)

t = - 6 ln(0.05)

t = 18 sec

Comment in case any doubt please rate my answer.....